Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
Cleva, Kefelian - 9:48 Wednesday 14 August 2019 (46663) Print this report
unlock

# Unlock at 6H49MN47SEC_UTC

I (Fred)  am responsible for  this unlock while we were in the atrium for locking the master laser frequency monitor (MLFM). Beside this light intervention I went at the rear of the master laser crate in order to plug a hole in the wooden box of the MLFM with some tape. IIt maybe that I have hit the fiber which bring the ML to the LB (stainless steel shield) although I pay attention to not hit it...

the plot show a raising oscillation on EOM_corr (4 kHz) which ends up in saturation and unlock

The signals ML_FREQ_I_MONIT / ML_FREQ_Q_MONIT refer to the ML frequency noise for frequency below 150 Hz and aboce 150 Hz, (see next entry for the releavant calibrations)

 

 

 

Images attached to this report
Search Help
×

Warning

×